write the equation of the line with slope 4/5 containing the point (3,7)
Answer by stanbon(75887)

write the equation of the line with slope 4/5 containing the point (3,7)
-----
Form: y = mx + b
----
You are told that y = 7 when x = 3 and you know m = 4/5
Solve for "b":
-------
7 = (4/5)3 + b
-----
7 = 12/5 + b
-----
35/5 = 12/5 + b
b = 23/5
--------
Equation:
y = (4/5)x + (23/5)
========================
Cheers,
Stan H.
----------